# Quantum Acoustics Demystifies the Strange Metals

### v1

@@ V.1.1 No sharp rise at low T
The $T^{4}$ rise in resistivity seen in normal metals below the Debye temperature will not apply, because the diffusion constant is already “pinned” at the Planckian rate as ever shorter modes awaken at higher temperature, making scant difference. Indeed, the $T^{4}$ rise is not seen in the strange metals.
@@ V.1.2 Mott-Ioffe-Regel bypass
This also explains why the Mott-Ioffe-Regel resistivity rollover is bypassed in the strange metals: the diffusion, responsible for the transport, is Planckian both before and after the MIR limit is passed, so the resistivity continues to rise as $T$.
@@ V.2 Electrons on a solid hydrogen surface
The experiment described now emboldened us to believe the Planckian diffusion we were seeing in WoW is truly ubiquitous.
Translated to the context of the strange metals, this experiment helps point the way to linear resistivity down to 0 Kelvin.
The experiment by Adams, and Adams and Paalanen [adams_localization_1987, adams_conductivity_1992], measures the mobility of nondegenerate electrons on a 2D solid hydrogen surface, perturbed by adsorbed Helium atoms, which release the electrons from bondage. Planckian diffusion $D\sim\hbar/m_{e}$ emerges, down to the lowest temperature measured, just below 2 K. The mobility is measured by magnetoresistance in a Corbino disk geometry (figure 10).
@@ V.4 Dynamical Planckian diffusion model
A simple, dynamic model is now defined, that obeys Planckian diffusion $D=\hbar/m^{*},$ leading (if any temperature is supposed, which is not necessary) to perfectly linear in temperature resistivity down to 0 Kelvin. The linearity in temperature comes from concentration gradients induced by applied fields, together with temperature-independent Planckian diffusion, not from any microscopic or mechanistic sensitivity of the dynamics to temperature. With the gradients established, Fickian diffusion reigns, causing transport, and linearity prevails.
@@ V.5 Drude peak displacement
The Drude peak in ordinary metals is a sharp rise in light absorption of the material at low frequency, resulting from nearly free carriers in the metal. At low frequencies the electrons are slowly dragged back and forth, as in a DC field; the absorption is basically due to DC resistivity. Drude-like scattering mechanistrange metals are at play. The Drude peak is absent in insulators or semiconductors without a high density of free electrons.
In a recent highlighted PRL [prlDDP], we addressed this question using WoW techniques.. See figure 13. The quantum-acoustical representation reveals a properly displaced Drude peak hiding in plain sight at higher temperatures within the venerable Fröhlich model: the optical conductivity obtained from thw WoW simulations the Kubo formalistrange metal exhibits a finite frequency maximum in the far-infrared range, while the near-d.c. conductivity is suppressed.
@@
Our WoW calculations were initiated with bound states of a frozen deformation potential, The deformation potential was then allowed to move normally and the conductivity calculated.s
@@ V.6 Polarons, CDW, and the linear approach to 0 Kelvin
More serious is the question of the local electron probability amplitude, since the effect of the electron locally on the lattice goes as $|\psi|^{2}$. However a recent paper, now published in PNAS,  [aydin_polaron_2025], allays these fears, showing how robust the numerical results are. Even the polaron formation dynamics is convincing: after a picosecond or so at temperatures under about 20 K, they form suddenly, although we do not yet understand just what happened the moment before. We can be fairly confident that a similar polaron formation scenario would apply in a full many-body implementation.
@@
Because polarons or CDW and nascent fluctuations are not phonons, we replace the idea of “electron-phonon” interactions, with “vibration-electron,” or better. “vibronic” interactions. The concept of vibronic interaction has been in use for decades in molecular physics and implies an electronic response to vibrational motion, and vise-versa, be it adiabatic or nonadiabatic. It maay be that the elusive understanding of the cuprates has not been in finding an emergent e-e fluid, but rather in developing a proper account of vibronic interactions, including polarons, CDW, and their associated fluctuations.
Typical $E_{d}$ in equation 7 is 10-30 eV. This is quite large, and may put us in good stead for the mean field approximations, which are proven to be exact regarding some inportant properties in the strong coupling limit [leopold2021derivation, leopold2021derivation, leopold2022landau].
In filling up the Fermi sea level by level, the total energy shift is the sum of individual level shifts up to the Fermi energy. This is worth noting because in a real sense, the electron-lattice coupling is strong at any temperature for the strange metals. For example, polarons, which are an extreme outcome of electron-lattice coupling, typically form below 50 K in our wave-on-wave simulations described below. They reflect strong vibronic coupling present even at 0K. The intrinsic electron-lattice coupling does not diminish with lower temperature. These are obvious enough comments, but if electron-phonon coupling is strong enough, and the Fermi energy and velocity are low enough, strange things can happen near 0 Kelvin. It is risky to assume the landscape that an electron sees near 0 K in the strange metals is serene, and incapable of scattering, just because the thermal part of the deformation potential has gone silent. The engine of Planckian diffusion leading to linear with temperature resistivity may be alive and well.
@@ V.7 Discussion and implications
One of the key features making the strange metals strange is perfectly linear in temperature resistivity over orders of magnitude in temperature, down even to perhaps 0.1 Kelvin, if superconductivity has been suppressed by a magnetic field. This holds in the strange metal phase of figure 8, and the linearity is seen in figure 9, though not down to 0.1 K. More than that, the slope of the resistivity vs. temperature is limited by a Planckian bound, in the form of a Planckian speed limit. This implies that quantum systems can diffuse not much faster or slower than a diffusion constant $D=\hbar/m^{*}$ allows.
Since we are in a speculative mode, we discuss a possible electron-lattice soup forming in the strange metal regime, even at low T, and dressing up electrons (or really holes) to act classically.
Electrons at the Fermi level in the cuprates carve backaction grooves in the deformation potential, much like a speedboat creates waves (see figure 15). We assume holes do this too. Supposing many speedboats on a lake with no wind, i.e. no thermal agitation, the waters are nonetheless anything but calm. The deformation potential is fluctuating, and acting back on the electrons, in equilibrium. Although the electron-lattice back action can cause CDW in other regimes, we do not need these, nor polarons, to explain the rough landscape that even cold temperatures present to the electrons. With such roiling, Planckian diffusion becomes plausible even at very low T. We suggest this as one possible explanation for even very low T linearity. Ths scenario has to occur in equilibrium, which means there is a kind of active electron-vibration soup formed, with electrons and holes in constant, two-way exchange of energy with the lattice. Electron energy would not be separable from the lattice in this strongly correlated soup. What is the compressibility of such a soup, as a function of T?
@@
We suspect that electrons are indeed strongly correlated in the strange metals, as is so often stated, but there is a big caveat: it is a side effect of strong vibronic interaction [strangemetal1], This happens through electron back-action on the lattice, regardless of the temperature. Since other electrons interact with the same vibronic soup, there are lattice mediated electron-electron interactions.
@@
The Planckian $D$ is not so much of a bound as it is ubiquitous; it can be violated in either direction, e.g. under extremely slow or extremely fast evolution of the medium. It seems clear that if the potential “scene” is locally and significantly altered before the time required to set up the delicate return-path coherence needed for localization, then localization will be ruined. That surely must lead to some kind of diffusion. Our advance is that it is Planckian diffusion - and a new “universality,” with the quotes allowing for exceptions.
@@
In the case of the electrons on solid hydrogen example, the disrupting agent is the presence of mobile helium atoms. Planckian diffusion in a moving random medium is the direct product, so to speak, of two universalities: the means of localization and the means of release.
@@
An important ramification of quantum diffusion is its isotropy, which was emphasized by Grissonnanche et. al. in 2021 [Grissonnanche2021], and taken as evidence against “hot spot” scenarios. Even if the deformation potential is anisotropic, for example if the band structure circular, the diffusion should still be isotropic. The reason is the universality we are now discussing: the diffusion, having neared a Planckian limit, is uncaring about details of the diffusing potential, and gives the same diffusion in any direction: no hot spots. Or put another way, the hot spots may be present but they fail to affect the diffusion very much. This “uncaring” property has colored many of the arguments against “phonons”, i.e. the lattice, when effects like isotope substitution did not produce the expected effect. We know, for example, that the deformation potential strength can vary many fold without much effect on the diffusion, which remains $D\sim\hbar/m$
@@ VI Summary
Planckian diffusion with $D=\hbar/m^{*}$ applies [ghost] widely; adding more scatterers doesn’t matter if the Planckian limit has already been reached. This also explains the lack of a $T^{4}$ rise of resistivity (as applies to 2d normal metals), as long as the Planckian regime has somehow been reached at low $T$. We argue that the landscape experienced by an electron is not flat, even at 0 K, in the strange metals.
Planckian diffusion is ubiquitous, and the diffusion constant $D=\hbar/m^{*}$ over huge parameter ranges remains no matter what agent is behind the stirring of the medium (but it must be stirred). Faster or more vigorous stirring has little effect on the Planckian rate. So mechanistrange metals can change with temperature, or doping, yet have almost exactly the same diffusion, and the same linear resistivity in T.
We again remark that this work started with the venerable workhorse Frölich Hamiltonian, and never had to deviate from those beginnings. We have shown that all the strange metal physics we found in contained with that model, and probably more will be found. The Hamiltonian itself stops at the lowest order electron-lattice strain expansion, yet when that interaction is allowed to act continuously and coherently, the treasure trove of physics it contains is revealed.

### v2

@@ V.1.1 No sharp rise at low T
The $T^{4}$ rise in resistivity seen in normal metals below the Debye temperature will not apply, because the diffusion constant is already “pinned” at the Planckian rate as ever shorter modes awaken at higher temperature, making scant difference. Indeed, the $T^{4}$ rise is not seen in the strange metal regime, or if present in parallel is overwhelmed by the fast rise T compared to T2 or T4 at low T.
@@ V.1.2 Mott-Ioffe-Regel bypass
The Mott-Ioffe-Regel resistivity rollover is bypassed in the strange metals for the following reason: the diffusion, responsible for the transport, is Planckian both before and after the MIR limit is passed, so the resistivity continues to rise as $T$.
@@ V.2 Electrons on a solid hydrogen surface
We now refer to the experiment by Adams, and Adams and Paalanen [adams_localization_1987, adams_conductivity_1992], measuring the mobility of nondegenerate electrons on a 2D solid hydrogen surface, perturbed by adsorbed Helium atoms. The Helium atoms release the electrons from bondage. Planckian diffusion $D\sim\hbar/m_{e}$ emerges, down to the lowest temperature measured, which is just below 2 K. The mobility is measured by magnetoresistance in a Corbino disk geometry (figure 10).
@@
This experiment emboldened us to believe the Planckian diffusion we were seeing in WoW is truly ubiquitous. It helps point the way to linear resistivity down to 0 Kelvin.
@@ V.4 Dynamical Planckian diffusion model
We present a simple, dynamic model that obeys Planckian diffusion $D\sim\hbar/m^{*},$ leading (if any temperature is supposed, which is not necessary) to perfectly linear in temperature resistivity down to 0 Kelvin. The linearity in temperature comes from concentration gradients induced by applied fields, together with temperature-independent Planckian diffusion, not from any microscopic or mechanistic sensitivity of the dynamics to temperature. With the gradients established, Fickian diffusion reigns, causing transport, and linearity prevails.
@@ V.5 Drude peak displacement
The Drude peak in ordinary metals is a sharp rise in light absorption of the material at low frequency, resulting from nearly free carriers in the metal. At low frequencies the electrons are slowly dragged back and forth, as in a DC field; the absorption is basically due to DC resistivity. The Drude peak is absent in insulators or semiconductors without a high density of free electrons.
Fratini and Ciuchi considered Drude peak displacement in this and similar contexts and assigned it to unspecified slow bosonic fluctuations[infrared]. We specifically modeled the exiled (see taboo above) lattice vibrations and found that they were responsible, after all:
In a recent highlighted PRL [prlDDP], we addressed this question using WoW techniques. See figure 13. The quantum-acoustical representation reveals a properly displaced Drude peak hiding in plain sight at higher temperatures within the venerable Fröhlich model: the optical conductivity obtained from thw WoW simulations the Kubo formalistrange metal exhibits a finite frequency maximum in the far-infrared range, while the near-d.c. conductivity is suppressed.
@@
Our WoW calculations were initiated with bound states of a frozen deformation potential, The deformation potential was then allowed to move normally and the conductivity was calculated numerically.
@@ V.6 Polarons, CDW, and the linear approach to 0 Kelvin
More serious is the question of the local electron probability amplitude, since the effect of the electron locally on the lattice goes as $|\psi|^{2}$. However a recent PNAS paper  [aydin_polaron_2025], allays these fears, showing how robust the numerical results are. Even the polaron formation dynamics is convincing: after a picosecond or so at temperatures under about 20 K, they form suddenly, although we do not yet understand just what happened the moment before. We can be fairly confident that a similar polaron formation scenario would apply in a full many-body implementation.
@@
Because polarons or CDW and nascent fluctuations are not phonons, we replace the idea of “electron-phonon” interactions, with “vibration-electron,” or better, “vibronic” interactions. The concept of vibronic interaction has been in use for decades in molecular physics and implies an electronic response to vibrational motion, and vise-versa, be it adiabatic or nonadiabatic. It maay be that the elusive understanding of the cuprates has not been in finding an emergent e-e fluid, but rather in developing a proper account of vibronic interactions, including polarons, CDW, and their associated fluctuations.
Typical $E_{d}$ in equation 7 is 10-30 eV. This is quite large and may put us in good stead for the mean field approximations, which are proven to be exact regarding some important properties in the strong coupling limit [leopold2021derivation, leopold2021derivation, leopold2022landau].
In filling up the Fermi sea level by level, the total energy shift is the sum of individual level shifts up to the Fermi energy. This is worth noting because in a real sense, the electron-lattice coupling is strong at any temperature for the strange metals. For example, polarons, which are an extreme outcome of electron-lattice coupling, typically form below 50 K in our wave-on-wave simulations described below. They reflect strong vibronic coupling present even at 0K. The intrinsic electron-lattice coupling does not diminish with lower temperature. These are obvious enough comments, but if electron-phonon coupling is strong enough, and the Fermi energy and velocity are low enough, strange things can happen near 0 Kelvin. It is risky to assume the landscape that an electron sees near 0 K in the strange metals is serene, and incapable of scattering, just because the thermal part of the deformation potential has gone nearly silent. The engine of Planckian diffusion leading to linear with temperature resistivity may be alive and well.
@@ V.7 Discussion and implications
One of the key features making the strange metals strange is perfectly linear in temperature resistivity over orders of magnitude in temperature, down even to perhaps 0.1 Kelvin, if superconductivity has been suppressed by a magnetic field. This holds in the strange metal phase of figure 8, and the linearity is seen in figure 9, though not down to 0.1 K. More than that, the slope of the resistivity vs. temperature is limited by a Planckian bound, in the form of a Planckian speed limit.
Since we are in a speculative mode, we discuss a possible electron-lattice soup forming in the strange metal regime, even at low T, and dressing up electrons (or really holes) to act nondegenerate.
Electrons at the Fermi level in the cuprates carve backaction grooves in the deformation potential, much like a speedboat creates waves (see figure 15). We assume holes do this too. Supposing many speedboats on a lake with no wind, i.e. no thermal agitation, the waters are nonetheless anything but calm. The deformation potential is fluctuating, and acting back on the electrons, in equilibrium. Although the electron-lattice back action can cause CDW in other regimes, we do not need these, nor polarons, to explain the rough landscape that even cold temperatures present to the electrons. With such roiling, Planckian diffusion becomes plausible even at very low T. We suggest this as one possible explanation for even very low T linearity. This scenario has to occur in equilibrium, which means there is a kind of active electron-vibration soup formed, with electrons and holes in constant, two-way exchange of energy with the lattice. Electron energy would not be separable from the lattice in this strongly correlated soup. What is the compressibility of such a soup, as a function of T?
@@
We suppose that electrons are indeed strongly correlated in the strange metals, as is so often stated, but there is a big caveat: it is an effect of strong vibronic interaction [strangemetal1], This happens through electron back-action on the lattice, regardless of the temperature. Since other electrons interact with the same vibronic soup, there are lattice mediated electron-electron interactions.
@@
The Planckian $D$ is not so much of a bound as it is ubiquitous; it can be violated in either direction, e.g. under extremely slow or extremely fast evolution of the medium. It seems clear that if the potential “scene” is locally and significantly altered before the time required to set up the delicate return-path coherence needed for localization, then localization will be compromised. That surely must lead to diffusion. Our advance is that it is Planckian diffusion - and a new “universality,” with the quotes allowing for exceptions.
@@
In the case of electrons on solid hydrogen, the disrupting agent is the presence of mobile helium atoms. Planckian diffusion in a moving random medium is the direct product, so to speak, of two universalities: the means of localization and the means of release.
@@
An important ramification of quantum diffusion is its isotropy, which was emphasized by Grissonnanche et. al. in 2021 [Grissonnanche2021], and taken as evidence against “hot spot” scenarios. Quoting, “we find … a T‐linear scattering rate that saturates the Planckian limit … Remarkably, we find that this Planckian scattering rate is isotropic, i.e. it is independent of direction, in contrast with expectations from ‘hot‐spot’ models.”
The reason for the isotropy is the universality we are now discussing: the diffusion, having neared a Planckian limit, is uncaring about details of the diffusing potential, and yields the same diffusion in any direction. Or put another way, the hot spots may be present but they fail to affect the diffusion very much. This “uncaring” property has colored many of the arguments against “phonons”, i.e. the lattice, when effects like isotope substitution did not produce the expected effect. We know, for example, that the deformation potential strength can vary many fold without much effect on the diffusion, which remains $D\sim\hbar/m$
@@ VI Summary
Planckian diffusion is ubiquitous, and the diffusion constant $D=\hbar/m^{*}$ over huge parameter ranges remains no matter what agent is behind the stirring of the medium (but it must be stirred). Faster or more vigorous stirring has little effect on the Planckian rate. So strange metals can change temperature, or doping, yet have almost exactly the same diffusion, and the same linear resistivity in T.
We again remark that this work started with the venerable workhorse Frölich Hamiltonian, and never had to deviate from those beginnings. We have shown that much strange metal physics can found in contained with that model, and probably more will be found. The Hamiltonian itself stops at the lowest order electron-lattice strain expansion, yet when that interaction is allowed to act continuously, coherently, and to infinite order, the treasure trove of physics it contains is revealed.